Diagonals of a rhombus are of lengths 16cm and 12cm.find the length of it's sides
Answer by Theo(13342)

all sides of a rhombus are equal in length.
the diagonals of the rhombus intersect at a 90 degree angle.
the diagonals and the sides of the rhombus form right triangles.
one leg of these right triangles is equal to 8 cm in length.
the other leg of these right triangles is equal to 6 cm in length
that would be half the length of each diagonal.
the sides of the triangle form the hypotenuse of these right triangles.
the formula is:
hypotenuse squared = one leg squared plus other leg squared.
this makes the hypotenuse squared equal to 8^2 + 6^2 = 64 + 36 = 100
the hypotenuse is the square root of 100 which makes the hypotenuse equal to 10.
the sides of the rhombus are equal to 10 cm.